Blossom Camping Supply sells marine-grade hatches. The hatches sell for $16 each. The cost of each hatch is $9. The only other costs are fixed costs of $11,998. During the current period, Blossom sold 1,920 units. What is the Contribution margin per unit? Contribution margin ratio? Breakeven point in units and in Dollars? Margin of safety in units and in dollars?
Answer #1

Selling price per unit = $16

Variable cost per unit = $9

Fixed cost = $11,998

Contribution margin per unit = Selling price per unit - Variable cost per unit

= 16-9

= $7

Contribution margin ratio = Contribution margin per unit/ Selling price per unit

= 7/16

= 43.75%

Break even point in units = Fixed cost/ Contribution margin per unit

= 11,998/7

= 1,714 units

Break even point in dollars= Fixed cost/ Contribution margin ratio

= 11,998/43.75%

= $27,424

Actual sales in units = 1,920

Actual sales in dollars = Selling price per unit x Number of units sold

= 16 x 1,920

= $30,720

Margin of safety in units = Actual sales in units - Break even sales in units

= 1,920-1,714

= 206 units

Margin of safety in dollars= Actual sales in dollars- Break even sales in dollars

= 30,720-27,424

= $3,296

Please give a positive rating if you are satisfied with this solution and if you have any query kindly ask.Thanks!!!